# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v1

@@ IV-B3 Denoising Diffusion Probabilistic Models
$+\sum_{t=2}^{T}\mathbb{E}_{q(\mathbf{y}_{t}|\mathbf{Y}_{0})}\operatorname{KL}[q(\mathbf{y}_{t-1}|\mathbf{y}_{t})||p_{\phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]$
@@
As denoted, the regularization term is assumed to be zero, since we assume that a sufficient amount of steps $T$ are taken such that $q(\mathbf{y}_{T}|\mathbf{y}_{0})$ is approximately normally distributed. With the reparameterization trick, the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$. As such, the KL-divergence in the second term can be interpreted as predicting $\mathbf{Y}_{0}$, the source noise $\bm{\epsilon}$ or the score $\nabla_{\mathbf{Y}_{t}}\log q(\mathbf{y}_{t})$ (gradient of the data log-likelihood) from $\mathbf{Y}_{t}$ depending on the parameterization, and is in almost all instances approximated with a U-Net [2].
It has also be proposed to model Bernoulli noise instead of Gaussian [93, 94, 95, 96]. However, most methodologies vary in the conditioning of the reverse process on the input image [97, 98, 99, 100]. For instance, Wolleb et al. [97] concatenate the input image with the noised segmentation mask. Wu et al. [98] insert encoded image features to the U-Net bottleneck. Additionally, information on predictions $\mathbf{Y}_{t}$ at a time step $t$ is also provided in the intermediate layers of the conditioning encoder. This is performed by applying the Fast Fourier Transform (FFT) on the U-Net encoding, followed by a learnable attentive map and the inverse FFT. The procedure of applying attention on the spectral domain of the U-Net encoding has also been done with transformers in follow-up work [99]. Segdiff [100] also encode both current time step and input image, but combine the extracted features by simple summation before applying the U-Net.
Figure 10: An illustration of the Diffusion Probabilistic Models [90]. The model learns to remove noise that has gradually been added to the input image.
@@ IV-C Test-time augmentation
An image $\mathbf{X}$ can be understood as a one-of-many visual representations of the object of interest. For example, systematic noise, translation or rotation result in many instances realistic variations that approximately retain image semantics. Hence, data augmentation [31] has been used at test time (hence the name test-time augmentation, or TTA) for image classification to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function [101]. This technique has been applied to image segmentation as well [102, 103, 104, 105]. By randomly augmenting input images with invertible transformation $T$ as $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$. Repeatedly performing this procedure results in a set segmentations masks, which can serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$.
@@ V Epistemic uncertainty
The crucial difference between epistemic and aleatoric uncertainty is that the former is related to model ignorance, while the latter reflects statistical ambiguity inherent in the data. Epistemic uncertainty can be further categorized into two distinct types [12]. The first type pertains uncertainty related to the capacity of the model. For example, under-parameterized models or approximate model posteriors can become too stringent to appropriately resemble the true posterior. The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, this is also known as model uncertainty. Nevertheless, given the complexity of contemporary parameter-intensive CNNs, the model uncertainty is often assumed to be negligible. A more significant contribution to the epistemic uncertainty is due to the limited data availability, known as approximation uncertainty, and can often be reduced by collecting more data. Both model and approximation uncertainty contribute to the epistemic uncertainty.
Unfortunately, evaluation of the true Bayesian posterior (formulated in Equation (1)) is inhibited by the intractability of the data-likelihood in the denominator. Hence, extensive efforts have been taken to obtain viable approximations such as using Mean-Field Variational Inference [8], Markov Chain Monte Carlo (MCMC) [106], Monte-Carlo Dropout [107, 108], Model Ensembling [109], Laplace approximations [110], Stochastic Gradient MCMCs [111, 112, 113], assumed density filtering [114] and expectation propagation [115, 116]. We refer to any neural network that approximates the Bayesian posterior of the model parameters as a Bayesian Neural Network (BNN). The following sections treat methodologies that have been applied within the context of this paper, which are usually straightforward extensions of BNN networks used for conventional regression and classification tasks. Additionally, an illustration of these techniques are presented in Figure 11.
@@ V-A Variational Inference
Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$, to approximate $p(\bm{\theta}|\mathbf{y},\mathbf{x})$. Then, we can achieve Variational Inference (VI) w.r.t. to the parameters by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior can be written as
@@
where the parameter-independent terms are constant and therefore excluded from notation. A popular choice for the approximated variational posterior is a the Gaussian distribution, i.e. a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary Gaussian prior densities. However, the priors can also be learned through Empirical Bayes [117]. Furthermore, backpropagation is possible with the reparameterization trick [10] and within this context, the procedure is referred to as Bayes by Backprop (BBB) [8]. During testing, a sample-based approach is utilized to approximate the posterior. Since using several parameter permutations effectively enriches the hypothesis space due to the model combining effect and can already improve performance [118, 119, 120].
@@ V-B Monte Carlo Dropout
Dropout, a common technique used to regularize neural networks [121], can mimic sampling from an implicit parameter distribution, $q(\tilde{\bm{\theta}}|\bm{\theta},p)$, defined as
$\mathbf{n}$ $\sim\operatorname{Bernoulli}(p)$ (35a)
@@
with probability $p$ and $\mathbf{n}$ operating element-wise on the parameters. Using Dropout can also be interpreted as a first-order equivalent $L_{2}$-regularization with additional transforming the input by the inverse diagonal Fisher information matrix [121]. With Monte-Carlo Dropout (MC Dropout), the random node switching is continued during testing, effectively sampling new sets of parameters. While seemingly arbitrary, it has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process [107]. In this manner, the authors showcase such method is able to provide multi-modal estimates of the model uncertainty.
As noted by Gal et al. [122], the model output variance is balanced with the weight magnitudes rather than the dropout rate $p$, which is usually optimized through grid search or simply fixed to $0.5$. Hence, the authors propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, such that uncertainty estimates are governed by $p$. As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution [123, 87], to enable path-wise derivatives through $p$.
@@ V-C Model Ensembling
As mentioned earlier, Monte Carlo dropout effectively optimizes over a set of sparse neural networks. This ensemble can also be designed in an explicit manner. Let us define the set of functions $\bm{f}=\{f^{n}_{\bm{\theta}}\}_{n=1}^{N}$, with $N$ representing the number of models in the ensemble. Then, it is relatively simple to obtain $\bm{\Theta}=\{\bm{\theta}_{n}\}_{n=1}^{N}$, which can be interpreted as samples from an approximate posterior. Ensembling in only the latter parts of a neural network (typically the decoder) is referred to as M-heads, i.e. the network has multiple outputs. Often, the $N$ obtained parameters are from $N$ separate training sessions. However, it has also been shown effective to ensemble from single training session by saving the parameters at multiple stages or training with different weight initializations [124, 125, 126].
@@
Figure 11: Varying techniques to sample parameters for the convolutional kernels. An approximation of the parameter density can be made (a), taking samples can be mimicked (b), or an ensemble of $N$ different configurations can explicitly be modeled (c).
A closely related concept to ensembling is known as Mixture of Experts (MoE), where each model in the ensemble (an ‘expert’) is trained on specific subsets of the data [127]. In such settings, a gating mechanism is usually applied after combining the expert hypotheses. While uncommon, incorporating all predictions can also be regarded as ensembling technique.
@@ VI Applications
This section explores literature that employs uncertainty-based downstream tasks on segmentation models. These include estimating the segmentation mask distribution subject to observer variability (Section VI-A), model introspection (i.e. ability to self-assess, Section VI-B), improved generalization (Section VI-C) and reduced labeling costs using Active Learning (Section VI-D).
@@ VI-A Observer variability
After observing sufficient data, the variability in the predictive distribution is often considered to be negligible and is therefore omitted. Nevertheless, this assumption becomes excessively strong in ambiguous modalities, where its consequence is often apparent with multiple varying, yet plausible annotations for a single image. Additionally, such annotations can also vary due to differences in expertise and experience of annotators. This phenomenon of inconsistent labels across annotators is known as the inter-observer variability, while variations from a single annotator is referred to as the intra-observer variability (see Figure 12).
To contextualize this phenomenon within the framework of uncertainty quantification, annotators can be treated as models themselves. For example, consider $K$ separate annotators modeled through parameters $\bm{\phi}_{k}$ with $k=1,2,...,K$. For a simple segmentation task, it can be expected that $\operatorname{Var}[p(\bm{\phi}_{k})]\rightarrow 0$. In other words, each annotator is consistent in their delineation and the intra-observer variability is low. For cases with consensus across experts, i.e. negligible inter-observer variability, the marginal converges to $\operatorname{Var}[p(\bm{\phi})]\rightarrow 0$. Asserting these two assumptions, it is valid to simply consider a point estimate of the posterior. Yet, this rarely the case in many real-life application and, as such, explicitly modeling the involved distributions becomes imperative.
Figure 12: A visualization of intra-observer variability in parameter space (left) and inter-observer variability in data space (right).
For evaluation, a common metric is to minimize the squared distance between arbitrary mean embeddings of the ground truth and predicted annotations using the kernel trick. This know as Maximum Mean Discrepancy (MMD) or the Generalized Energy Distance (GED) [128], denoted as
$\operatorname{GED}^{2}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]$
@@
with marginals $P_{\mathbf{Y}}$ and $P_{\hat{\mathbf{Y}}}$, representing the true and predictive segmentation distribution, and some kernel $k:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, usually the 1--IoU or 1--Dice score. An alternative metric known as Hungarian Matching (HM) compares the predictions against the ground-truth labels through a cost matrix [71]. Subsequently, the unique optimal coupling between the two sets that minimizes the average cost is determined through combinatorial optimization algorithm. This can also be formally denoted as finding the permutation matrix $\mathbf{P}$ subject to the objective
$\operatorname{HM}(Y,\hat{Y})=\min_{\mathbf{P}}\frac{1}{N^{2}}\operatorname{Tr}(\mathbf{P}\mathbf{M}),$ (37)
where $M_{i,j}=k(y_{i},\hat{y}_{j})$ and $N^{2}$ represents the number of elements of the matrix.
For accurate comparison and evaluation of the problem, benchmarking is often constrained to publicly available multi-annotated data. For example, the Lung Image Database Consortium and Image Database Resource Initiative (LIDC-IDRI) [129] contains manually annotated lesions from lung patients and is one of the few datasets that has extensively been evaluated with appropriate metrics (see Table I). Multiple versions have been used, either with 15,096 (LIDCv1) or 12,816 (LIDCv2) patches. LIDCv1 is employed with a 60:20:20 split while for LIDCv2 this is 70:15:15 for train:validation:test. LIDCv2 has also been used with threefold cross validation (LIDCv2-cv), with either a 90:10 or 80:20 train:test split.
Other less commonly datasets include the CityScapes [130] dataset, which contains street view images of German cities from the perspective of a driving car, have been used to artificially create class-level label ambiguity [71, 58, 78]. Some classes are switched to an arbitrary auxiliary class with some probability $p$. Since the underlying probabilities are known, the empirical fraction of the classes from the model predictions can directly be compared to the ground-truth values of $p$. Furthermore, the QUBIQ Challenge [131] contains MRI and CT data from varying organs. Also, the retinal fundus images for glaucoma analysis (RIGA) [132] dataset contains delineations of optic cup and disc boundaries by six experienced ophthalmologists.
@@
Method Year GED16 HM-IoU16 GED16 HM-IoU16
PU-Net [58] 2018 $0.310\pm 0.010$ $0.552\pm 0.000$ $0.320\pm 0.030$ $0.500\pm 0.030$
HPU-Net [71] 2019 $0.270\pm 0.010$ $0.530\pm 0.010$ $0.270\pm 0.010$ $0.530\pm 0.010$
PhiSeg [72] 2019 $0.262\pm 0.000$ $0.586\pm 0.000$
SSN [50] 2020 $0.259\pm 0.000$ $0.558\pm 0.000$
CAR [56] 2021 - - $0.264\pm 0.002$ $0.592\pm 0.005$
JProb. U-Net [133] 2022 - - $0.262\pm 0.000$ $0.585\pm 0.000$
PixelSeg [53] 2022 $0.243\pm 0.010$ $0.614\pm 0.000$ $0.260\pm 0.000$ $0.587\pm 0.010$
MoSE [78] 2022 $0.218\pm 0.001$ $0.624\pm 0.004$
AB [134] 2022 $0.213\pm 0.001$ $0.614\pm 0.001$
CIMD [95] 2023 $0.234\pm 0.005$ $0.587\pm 0.001$
CCDM [94] 2023 $0.212\pm 0.002$ $0.623\pm 0.002$ $0.239\pm 0.003$ $0.598\pm 0.001$
Explicitly modeling the annotator distribution has been explored with an MoE approach (Section V-C), using data with consistent image-annotations pairs, provided that annotators have an intrinsically associated expertise [135]. Nevertheless, relying on the model to infer ambiguity in the parameters by observing the data can become quite burdensome. Hence, it can be much simpler to directly model the empirical stochasticity in the annotations and has been extensively explored in modalities such as Lung nodule detection in 2D [58, 72, 71, 136, 137, 70, 66, 68, 65, 50, 64], as well as 3D [76, 138, 139], Brain Tumor [140, 50, 97, 93, 95], White Matter Hyperintensities [141], Pulmonary Tumour Growth [142], Prostate [143, 73] and vascular [144], street scene [145, 71, 94], aerial imaging [100], optic cup [98, 146], abdominal multi-organ [99] and nuclei microscopy [96] segmentation. The overwhelming majority of research for this particular application has been executed with variants of conditional VAEs [58, 66, 64, 65, 70, 68, 71, 72, 137, 142, 147, 140, 148, 139, 144, 76, 73, 143, 138]. More recently, the growing popularity of DDPMs is also apparent in the field [94, 100, 97, 93, 95]. Also, GAN-based approaches have been also employed [56].

### v3

@@ IV-B3 Denoising Diffusion Probabilistic Models (DDPMs)
$+\sum_{t=2}^{T}\mathbb{E}_{q\left(\mathbf{y}_{t}|\mathbf{Y}_{0}\right)}\big{[}\operatorname{KL}[q(\mathbf{y}_{t-1}|\mathbf{y}_{t})\,||\,p_{\phi}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]\,\big{]}$
@@
As denoted, the regularization term is assumed to be zero, since it is assumed that a sufficient amount of steps $T$ are taken such that $q(\mathbf{y}_{T}|\mathbf{y}_{0})$ is approximately normally distributed. With the reparameterization trick [10], the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$. As such, the KL divergence in the second term can be interpreted as predicting $\mathbf{Y}_{0}$, the source noise $\bm{\epsilon}$ or the score $\nabla_{\mathbf{Y}_{t}}\log q(\mathbf{y}_{t})$ (gradient of the data log-likelihood) from $\mathbf{Y}_{t}$, depending on the parameterization. This is in almost all instances approximated with a U-Net [2].
Also, it has been proposed to model Bernoulli noise instead of Gaussian [94, 95, 96, 97]. However, most methodologies vary in the conditioning of the reverse process on the input image [98, 99, 100, 101]. For instance, Wolleb et al. [98] concatenate the input image with the noised segmentation mask. Wu et al. [99] insert encoded image features to the U-Net bottleneck. Additionally, information on predictions $\mathbf{Y}_{t}$ at a time step $t$ is provided in the intermediate layers of the conditioning encoder. This is performed by applying the Fast Fourier Transform (FFT) on the U-Net encoding, followed by a learnable attention map and the inverse FFT. The procedure of applying attention on the spectral domain of the U-Net encoding has also been implemented with transformers in follow-up work [100]. Segdiff [101] also encode both current time step and input image, but combine the extracted features by simple summation prior to inferring it through the U-Net.
Figure 10: Illustration of the Diffusion Probabilistic Models [90]. The model learns to remove noise that has been gradually added to the input image.
@@ IV-C Test-time augmentation
An image $\mathbf{X}$ can be understood as a one-of-many visual representations of the object of interest. For example, systematic noise, translation or rotation result in many realistic variations that approximately retain image semantics. Hence, data augmentation [31] has been used at test time (explaining the name test-time augmentation, or TTA) for image classification, to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function [102]. This technique has been applied to image segmentation as well [103, 104, 105, 106]. By randomly augmenting input images with invertible transformation $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$. Repeatedly performing this procedure results in a set of segmentation masks, which serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$.
@@ V Epistemic uncertainty
The crucial difference between epistemic and aleatoric uncertainty is that the former relates to model ignorance, while the latter reflects statistical ambiguity inherent in the data. Epistemic uncertainty can be further categorized into two distinct types [12]. The first type pertains uncertainty related to the capacity of the model. For example, under-parameterized models or approximate model posteriors, that can become too stringent to appropriately resemble the true posterior. The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, which is also known as model uncertainty. Nevertheless, given the complexity of contemporary parameter-intensive CNNs, the model uncertainty is often assumed to be negligible. A more significant contribution to the epistemic uncertainty is due to the limited data availability, known as approximation uncertainty, and can often be reduced by collecting more data. Both model and approximation uncertainty contribute to the epistemic uncertainty.
Unfortunately, evaluation of the true Bayesian posterior (formulated in Equation (1)) is inhibited by the intractability of the data-likelihood in the denominator. Hence, extensive efforts have been taken to obtain viable approximations, such as using Mean-Field Variational Inference [8], Markov Chain Monte Carlo (MCMC) [107], Monte-Carlo Dropout [108, 109], Model Ensembling [110], Laplace approximations [111], Stochastic Gradient MCMCs [112, 113, 114], assumed density filtering [115] and expectation propagation [116, 117]. We refer to any neural network that approximates the Bayesian posterior of the model parameters as a Bayesian Neural Network (BNN). The following sections treat applied methodologies to segmentation, which are usually straightforward extensions of conventional regression and classification BNN networks. These techniques are illustrated in Figure 11.
@@ V-A Variational Inference
Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$, to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$. Then, we can employ Variational Inference (VI) w.r.t. to the parameters, by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior as
@@
where the parameter-independent terms are constant and therefore excluded from notation. A popular choice for the approximated variational posterior is the Gaussian distribution, i.e. a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary Gaussian prior densities. However, the priors can be also learned through Empirical Bayes [118]. Furthermore, backpropagation is possible with the reparameterization trick [10] and within this context, the procedure is referred to as Bayes by Backprop (BBB) [8]. During testing, a sample-based approach is utilized to approximate the posterior. Using several parameter permutations effectively enriches the hypothesis space due to the model combining effect, and can improve performance [119, 120, 121].
@@ V-B Monte Carlo Dropout
Dropout is a common technique used to regularize neural networks [122] and can mimic sampling from an implicit parameter distribution $q(\tilde{\bm{\theta}}|\bm{\theta},p)$. This implicit distribution is defined as
$\mathbf{n}$ $\sim\operatorname{Bernoulli}(p),$ (34a)
@@
with probability $p$ and vector $\mathbf{n}$ operating element-wise on the parameters. Using Dropout can also be interpreted as a first-order equivalent $L_{2}$ regularization with additionally transforming the input with the inverse diagonal Fisher information matrix [122]. Using Monte-Carlo Dropout (MC Dropout), the random node switching is continued during test time, effectively sampling new sets of parameters. While seemingly arbitrary, it has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process [108]. In this manner, such method is able to provide multi-modal estimates of the model uncertainty.
As noted by Gal et al. [123], the model output variance is balanced with the weight magnitudes rather than the dropout rate $p$, which is usually optimized through grid search or simply set to $0.5$. Hence, the authors propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, such that uncertainty estimates are governed by $p$. As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution [124, 87], to enable path-wise derivatives through $p$.
@@ V-C Model Ensembling
As mentioned earlier, MC dropout effectively optimizes over a set of sparse neural networks. Also, this ensemble can be designed in an explicit manner. To this end, we define the set of functions $\bm{f}=\{f^{n}_{\bm{\theta}}\}_{n=1}^{N}$, with $N$ representing the number of models in the ensemble. Then, it is relatively simple to obtain $\bm{\Theta}=\{\bm{\theta}_{n}\}_{n=1}^{N}$, which can be interpreted as samples from an approximate posterior. Ensembling in only the latter parts of a neural network (typically the decoder) is referred to as M-heads, i.e. the network has multiple outputs. Often, the $N$ obtained parameters are from $N$ separate training sessions. However, it has also been proven effective to ensemble from a single training session, by saving the parameters at multiple stages or training with different weight initializations [125, 126, 127].
@@
Figure 11: Three types of techniques are visualized to sample parameters of the convolutional kernels, (a) an approximation of the parameter density can be made, (b) taking samples can be mimicked with dropout during test time, or, (c) an ensemble of $N$ different configurations can be explicitly modeled.
A closely related concept to ensembling is known as Mixture of Experts (MoE), where each model in the ensemble (an ‘expert’) is trained on specific subsets of the data [128]. In such settings, a gating mechanism is usually applied after combining the expert hypotheses. While uncommon, incorporating all predictions can also be regarded as an ensembling technique.
@@ VI Applications
This section explores literature that employs uncertainty-based downstream tasks on segmentation models. These include estimating the segmentation mask distribution subject to observer variability (Section VI-A), model introspection (i.e. ability to self-assess, Section VI-B), improved generalization (Section VI-C), and reduced labeling costs using Active Learning (Section VI-D).
@@ VI-A Observer variability
After observing sufficient data, the variability in the predictive distribution is often considered to be negligible and is therefore omitted. Nevertheless, this assumption becomes excessively strong in ambiguous modalities, where its consequence is often apparent with multiple varying, yet plausible annotations for a single image. Additionally, such annotations can also vary due to differences in expertise and experience of annotators. The phenomenon of inconsistent labels across annotators is known as the inter-observer variability, while variations from a single annotator are referred to as the intra-observer variability (see Figure 12).
To contextualize this phenomenon within the framework of uncertainty quantification, annotators can be regarded as models themselves. For example, consider $K$ separate annotators modeled through parameters $\bm{\phi}_{k}$ with $k=1,2,...,K$. For a simple segmentation task, it can be expected that $\text{Var}[\,p(\bm{\phi}_{k})\,]\rightarrow 0$. In other words, each annotator is consistent in his delineation and the intra-observer variability is low. For cases with consensus across experts, i.e. yielding negligible inter-observer variability, the marginal approaches to $\text{Var}[\,p(\bm{\phi})\,]\rightarrow 0$. Asserting these two assumptions, it is valid to simply consider a point estimate of the posterior. Yet, this is rarely the case in many real-life applications and, as such, explicitly modeling the involved distributions becomes necessary.
Figure 12: Visualization of the intra-observer variability in parameter space (left) and the inter-observer variability in data space (right).
For evaluation, a commonly used metric minimizes the squared distance between arbitrary mean embeddings of the ground truth and predicted annotations using the kernel trick [129]. This metric is known as Maximum Mean Discrepancy (MMD) or the Generalized Energy Distance (GED) [130], denoted as
$\text{GED}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]$
@@
with marginals $P_{\mathbf{Y}}$ and $P_{\hat{\mathbf{Y}}}$, representing the true and predictive segmentation distribution, and a distance represented by kernel $k:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, usually the (1--IoU) or (1--Dice) score (squared by the GED). An alternative metric known as Hungarian Matching (HM) compares the predictions against the ground-truth labels through a cost matrix [74]. Subsequently, the unique optimal coupling between the two sets that minimize the average cost is determined through a combinatorial optimization algorithm. This can be also formally denoted as finding the permutation matrix $\mathbf{P}$, subject to the objective
$\text{HM}(\mathcal{Y},\mathcal{\hat{Y}})=\frac{1}{N^{2}}\min_{\mathbf{P}}\text{Tr}(\mathbf{P}\mathbf{M}),$ (36)
where the elements of $\mathbf{M}$ are $M_{i,j}=k(\mathcal{Y}_{i},\mathcal{\hat{Y}}_{j})$, and $N^{2}$ represents the number of elements within the matrix.
For accurate comparison and evaluation of the problem, benchmarking is often constrained to publicly available multi-annotated data. For example, the Lung Image Database Consortium and Image Database Resource Initiative (LIDC-IDRI) [131] contains manually annotated lesions from lung patients and is one of the few datasets that has extensively been evaluated with appropriate metrics (see Table I). Multiple versions have been used, either with 15,096 (LIDCv1) or 12,816 (LIDCv2) patches. LIDCv1 is employed with a 60:20:20 split, while for LIDCv2 this is 70:15:15 for train:validation:test. LIDCv2 has also been used with threefold cross-validation (LIDCv2-cv), with either a 90:10 or 80:20 train:test split [63, 59, 73].
Other less commonly employed datasets include the CityScapes [132] dataset, which contains street-view images of German cities from the perspective of a driving car, have been used to artificially create class-level label ambiguity [74, 69, 81]. Classes are switched to an arbitrary auxiliary class with some probability $p$. Since the underlying probabilities are known, the empirical fraction of the classes from the model predictions can be directly compared to the ground-truth values of $p$. Furthermore, the QUBIQ Challenge [133] contains MRI and CT data from varying organs. Also, the retinal fundus images for glaucoma analysis (RIGA) [134] dataset contains delineations of optic cup and disc boundaries by 6 experienced ophthalmologists.
@@
Method GED16 HM-IoU16 GED16 HM-IoU16
PU-Net [69], 2018 $0.310\pm 0.010$ $0.552\pm 0.000$ $0.320\pm 0.030$ $0.500\pm 0.030$
HPU-Net [74], 2019 $0.270\pm 0.010$ $0.530\pm 0.010$ $0.270\pm 0.010$ $0.530\pm 0.010$
PhiSeg [75], 2019 $0.262\pm 0.000$ $0.586\pm 0.000$
SSN [50], 2020 $0.259\pm 0.000$ $0.558\pm 0.000$
CAR [56], 2021 - - $0.264\pm 0.002$ $0.592\pm 0.005$
JProb. U-Net [135], 2022 - - $0.262\pm 0.000$ $0.585\pm 0.000$
PixelSeg [53], 2022 $0.243\pm 0.010$ $0.614\pm 0.000$ $0.260\pm 0.000$ $0.587\pm 0.010$
MoSE [81], 2022 $0.218\pm 0.001$ $0.624\pm 0.004$
AB [136], 2022 $0.213\pm 0.001$ $0.614\pm 0.001$
CIMD [96], 2023 $0.234\pm 0.005$ $0.587\pm 0.001$
CCDM [95], 2023 $0.212\pm 0.002$ $0.623\pm 0.002$ $0.239\pm 0.003$ $0.598\pm 0.001$
Explicitly modeling the annotator distribution has been explored with an MoE approach (Section V-C), using data with consistent image-annotation pairs, provided that annotators have an intrinsically associated expertise [137]. Nevertheless, relying on the model to infer ambiguity in the parameters by observing the data, can become quite burdensome. Hence, it can be much simpler to directly model the empirical stochasticity in the annotations, which has been extensively explored in modalities such as Lung nodule detection in 2D [69, 75, 74, 138, 139, 68, 60, 62, 73, 50, 59], as well as 3D [79, 140, 141], Brain Tumor [142, 50, 98, 94, 96], White Matter Hyperintensities [143], Pulmonary Tumour Growth [144], Prostate [145, 76] and vascular [146], street scene [147, 74, 95], aerial imaging [101], optic cup [99, 148], abdominal multi-organ [100] and nuclei microscopy [97] segmentation. The overwhelming majority of research for this particular application has been executed with variants of conditional VAEs [69, 60, 59, 73, 68, 62, 74, 75, 139, 144, 149, 142, 150, 141, 146, 79, 76, 145, 140]. More recently, the growing popularity of DDPMs is apparent in the field [95, 101, 98, 94, 96]. Also, GAN-based approaches have been employed [56].